OpenStudy (anonymous):

Marie mixes 16 liters of 18% acid solution with a 27% acid solution, which results in 21% acid solution. Let x represent the number of liters of 27% solution used and let y represent the number of liters of mixture. Which system of equations can be used to find the number of liters of 27% acid solution in the mixture?

OpenStudy (anonymous):

A: y = x + 16; 0.27y = 0.21x + 2.88 B: y = x + 16; 0.21y = 0.27x + 2.88 C: y = x + 16; 0.21y = 0.27x + 4.32 D: y = x + 16; 0.27y + 2.88x = 0.21